OpenStudy (anonymous):

the quadratic equation whose roots are 1-3^1/2 and 1+3^1/2 is

OpenStudy (ahsome):

You are missing \(x\) somewhere

OpenStudy (ahsome):

@gilbert936, we are missing \(x\) somehwere

OpenStudy (anonymous):

i dont understand you

OpenStudy (anonymous):

you know what! can you solve it or not

OpenStudy (ahsome):

Ok. A quadratic equation has an \(x^2\) inside it. In your two roots, there is no \(x\), therefore, you cannot get a quadratic equation.

OpenStudy (ahsome):

In your two roots, there HAS to be an \(x\) in the roots

OpenStudy (ahsome):

Or else it would just equal a number. Tell me where the \(x\) is, then I can help solve it.

OpenStudy (anonymous):

you can only solve the equation using sum of roots and product of roots, try it you will surely get the answer and thanks for trying my question

OpenStudy (ahsome):

@gilbert936. We can do that, its just that you will get a number. Try it yourself: \[1-3^{1/2} *1+3^{1/2}\]\[=(1-3^{1/2} )(1+3^{1/2})\] Use the rule: \((a-b)(a+b)=a^2-b^2\)\[(1)^2-(3^{1/2})^2\]\[=1-3\]\[=2\] The problem is, we don't get a quadratic equation. A quadratic equation has a \(x^2\) inside of it. Since we don't have a \(2\) in either of the roots, there is no Quadratic equation, just a number

OpenStudy (anonymous):

your answer is wrong

OpenStudy (anonymous):

given alpha, beta as the roots of a quafratic equation, the equation is given as X^2-(alpha,beta)X + alphabeta=0 therefore alpha=1-13^1/2, beta=1+13^1/2 alpha+beta=2 alphabeta=(1-13^1/2)(1+13^1/2)=1-13=-12 the equation is X^2-2X-12